# EnScale: Temporally-consistent multivariate generative downscaling via proper scoring rules

### v2

@@ 5.4.4 Extremes
Table 4: Extreme evaluation regarding marginal quantiles as well as miscalibrationa
Method tas pr
AE $Q^{0.05}_{\text{Wi}}\downarrow$ AE $Q^{0.95}_{\text{Su}}\downarrow$ MCB $Q^{0.1}$ $\downarrow$ MCB $Q^{0.9}$ $\downarrow$ AE $Q^{0.05}_{\text{Wi}}\downarrow$ AE $Q^{0.95}_{\text{Su}}\downarrow$ MCB $Q^{0.1}$ $\downarrow$ MCB $Q^{0.9}$ $\downarrow$
NN-det 1.021 0.948 nan nan 0.007 6.960 nan nan
EasyUQ 0.539 0.479 0.060 0.064 0.000 2.459 0.028 0.039
Analogues 0.847 0.485 0.103 0.054 0.000 1.449 0.068 0.085
GAN 1.606 1.545 0.097 0.063 0.012 4.395 0.274 0.038
CorrDiff 1.069 1.086 0.056 0.050 0.044 5.327 0.088 0.047
EnScale 0.488 0.478 0.021 0.021 0.000 2.813 0.045 0.018
EnScale-t 0.673 0.817 0.025 0.018 0.000 2.869 0.047 0.019
Method sfcWind rsds
AE $Q^{0.05}_{\text{Wi}}\downarrow$ AE $Q^{0.95}_{\text{Su}}\downarrow$ MCB $Q^{0.1}$ $\downarrow$ MCB $Q^{0.9}$ $\downarrow$ AE $Q^{0.05}_{\text{Wi}}\downarrow$ AE $Q^{0.95}_{\text{Su}}\downarrow$ MCB $Q^{0.1}$ $\downarrow$ MCB $Q^{0.9}$ $\downarrow$
NN-det 0.457 0.929 nan nan 6.744 15.976 nan nan
EasyUQ 0.108 0.211 0.041 0.049 1.333 3.178 0.023 0.038
Analogues 0.102 0.254 0.076 0.095 1.110 3.730 0.064 0.119
GAN 0.279 0.775 0.064 0.068 7.620 13.761 0.091 0.038
CorrDiff 0.282 0.491 0.059 0.043 3.219 9.879 0.051 0.097
EnScale 0.118 0.278 0.013 0.015 2.623 5.488 0.017 0.024
EnScale-t 0.124 0.266 0.013 0.025 2.377 5.597 0.015 0.023
We focus on the capability of EnScale and the benchmark to capture extreme events. Firstly, we assess locationwise quantiles in Tab. 4. EnScale performs well compared to the benchmarks, but best scores are often reached by the analogues and EasyUQ. We note, however, that these estimates are conservative – per season, there are approximately $1000$ days in the test set, leading to estimation errors for extreme quantiles. Notably, the GAN and CorrDiff struggle to reproduce extreme quantiles accurately. In particular, they miss to predict the zero precipitation events correctly. The true 5 % -quantile in winter is zero, but the ML approaches do not capture this. EnScale successfully predicts zero rain events. This is an advantage of our data transformation (Sec. 3.2).
@@
Figure 9: Extreme quantiles for the summer season (June, July, August) compared to ALADIN63 driven by CNRM-CM5. The first column shows the 95 % - quantile of the RCM calculated separately in each location. The remaining columns show the error for different methods, calculated as the signed difference of quantiles in the samples minus the RCM’s quantiles. Additionally, the last column provides an assessment of “model uncertainty” where the difference of ALADIN63 to CCLM4-8-17 (both driven by CNRM-CM5) is shown. Similar results are obtained for pairs other than ALADIN63 driven by CNRM-CM5.
@@
A more detailed analysis of the 95 % - quantile in summer is shown in Fig. 9. This illustration reveals the direction of error, which varies across locations and variables. For example, EnScale overestimates extreme temperatures but underestimates extreme solar radiation in the majority of locations. For precipitation, absolute errors are highest in the Alpine region, and for surface wind, errors are largest over the Sea. Again, EnScale performs well in comparison to the benchmarks, with errors of similar magnitude as for the analogues and often lower than for GAN and CorrDiff.
@@
In order to provide an additional perspective on these results, we compare to uncertainties across different RCMs. As an example, we show differences between the quantiles of two RCMs, both driven by the same GCM. We find that the mismatches between RCMs are of similar magnitude to the errors of the ML models. Importantly, in most cases, disagreements between the RCMs exceed the errors of EnScale. This suggests that, despite some limitations, the emulators typically produce results that fall within the range spanned by multiple RCMs.
@@
Next, we consider calibration of low / high quantiles of the predicted conditional distribution (see “MCB Q” columns in Tab. 4 and Sec. 4.2.4 for details). Here, EnScale and EnScale-t usually reach the lowest errors, with the exact ordering of these three methods depending on the variable and quantile considered. This again exemplifies that EnScale is sometimes inferior to the analogues with respect to metrics evaluating the marginal distributions $p_{Y}$. For the conditional $p_{Y|X}$, however, EnScale is superior.
5.4.5 Temporal structure
Table 5: Error in lag-1 autocorrelations (ACF) for each variablea
Method tas pr sfcWind rsds
NN-det 0.01 0.17 0.08 0.11
EasyUQ -0.03 -0.07 -0.15 -0.03
Analogues -0.10 -0.12 -0.28 -0.04
GAN -0.06 -0.1 -0.21 0.02
CorrDiff 0.01 0.23 0.10 0.08
EnScale -0.08 -0.09 -0.22 -0.06
EnScale-t -0.01 -0.01 -0.04 -0.01
aWe calculate ACFs in each location for both the samples and the RCM data.We average the signed difference of samples’ ACF minus target ACF over all locations.
We investigate the temporal structure of the generated samples, that is, how accurately the methods capture correlation between consecutive time steps. To this end, Tab. 5 shows differences in lag-1 (i.e., a single time step) autocorrelations (ACFs) between the samples and the RCM data.
We find that all methods except NN-det and CorrDiff have negative differences, i.e., they underestimate the local autocorrelation. CorrDiff accurately reproduces autocorrelations for temperature, but overestimates them for all other variables (see App. A.2 for more details). EnScale-t is able to reduce the error substantially and has only a small tendency to underestimate the autocorrelation for sfcWind. To ensure that the error reduction does not stem from canceling errors due to opposite signs, we compare absolute errors in Tab. A.2. We find that EnScale-t is superior to the others and achieves the lowest deviation from the true ACFs across all variables.
Capturing autocorrelations is particularly relevant when analyzing statistics aggregated over multiple days, such as in 5-day heatwaves. Underestimating autocorrelation can lead to an underrepresentation of variability in multi-day averages, therefore potentially underestimating risk. These findings point to limitations of existing ML-based downscaling approaches that fail to include temporal correlation on a local scale. EnScale-t has small room for improvement, especially for wind over the Alps (not shown). This is likely due to the approximation of modeling temporal consistency only on the level of the coarsened RCM (see Sec. 2.3). Nevertheless, even with this approximation, we can see a clear benefit compared to the other models.
Figure 10: As Fig. 6, omitting CorrDiff, but presenting all four target variables instead.
We show an exemplary time series in Fig. 10 for the spatial mean over one year. We compare EnScale, EnScale-t, and the underlying RCM. Most importantly, we find that generated time series by both, EnScale and EnScale-t, closely match the ground truth over the entire year and capture the seasonal cycle. On most days, the truth lies within the range of the 9 samples.
The illustration highlights the stability of the time series of the autoregressive rollout for our temporally-consistent coarse model. Both the individual time series as well as the uncertainty range remain stable over time. We also tested longer rollout of our generated time series and found that they are stable even for several years (not shown). Both the time series as well as the loss do not increase over time. The stability of rollout is an open discussion in the case of ML-based weather forecasting [Price2025ProbabilisticLearning, Karlbauer2024AdvancingMesh]. In our case, the additional input from the GCM likely helps to stabilize the rollout and prevent potential drifts.
The insets in the plots highlight qualitative differences between generated time series by EnScale-t and EnScale. The time series for the EnScale-t tend to be smoother because they take into account the temporal structure, especially for temperature. The time series by EnScale is not necessarily always consistent over time as each day is generated independently of the previous. The uncertainty bands are similar for both approaches. In times where temporal autocorrelation is high (especially visible for temperature), the uncertainty range for the EnScale-t is larger because including temporal structure allows us to sample a larger variety of possible trajectories.
@@ 5.5 Variable dependencies
Table 6: Multi-variable evaluationa
Method Correlations Miscalibration
pr-sfcWind sfcWind-rsds tas-pr pr-sfcWind sfcWind-rsds tas-pr
NN-det 0.125 0.062 0.060 nan nan nan
EasyUQ 0.166 0.217 0.055 0.563 0.790 0.389
Analogues 0.026 0.029 0.025 0.355 0.331 0.325
GAN 0.063 0.056 0.042 0.098 0.092 0.123
CorrDiff 0.126 0.094 0.117 0.195 0.192 0.213
EnScale 0.044 0.029 0.031 0.055 0.054 0.067
EnScale-t 0.042 0.031 0.033 0.056 0.054 0.067
@@
So far, our evaluation considered each target variable separately. Next, we evaluate the capability of our model and the benchmarks to capture the full distribution jointly for the four targets in Tab. 6.
Firstly, we consider pairwise correlations between variables for three selected variable pairs. For all pairs, the analogue benchmark reaches best or second-best scores. This is because the analogues only reproduce existing RCM data. By construction, the dependencies between variables are correct for the existing data. Results by EnScale are similar. In contrast, EasyUQ which models each variable independently of others, shows much higher errors. This is another example for the benefits of our multivariate approach.
Figure 11: Correlations between pairs of variables. The first column shows the RCM’s pairwise correlation between the variables in each location. The other columns show the error of EnScale and other methods, i.e., the signed difference of correlations in the samples minus the RCM. This figure is based on the results for CNRM-CM5 and ALADIN63 only, but other pairs reveal a similar pattern. The last column “model uncertainty” refers to the difference in correlations of two example RCMs, ALADIN63 and CCLM4-8-17, both driven by CNRM-CM5.
Fig. 11 presents the pairwise correlations in more detail. The patterns for the RCM show local details, especially in mountain regions. Errors for EnScale and EnScale-t are small, comparing favorably to the benchmarks. Sign and magnitude of errors vary across space. As in Sec. 5.4.4, we again put ML model errors into context to RCM disagreements. We present differences between the inter-variable correlations of two RCMs. Our findings are similar to what was observed previously: RCM differences are about the same magnitude as ML model errors, with EnScale often reaching smaller errors than the RCM mismatches.
The next metrics evaluates dependencies of the variables conditional on GCM data $X$: we consider calibration of the pairwise difference of variables (last three columns of Tab. 6, see Sec. 4.2.6 for details). Now, performance of the analogues is worse. Additional evaluation shows that they are again underdispersed (not shown). Instead, EnScale or EnScale-t achieve the best results. This finding again highlights the difference when analyzing marginals vs. conditional distributions, as in Sec. 5.4.4.
@@
We note that the above metrics only evaluate pairwise distributions between variables. Assessing the full joint distribution is not considered. In addition, we don’t explicitly evaluate physical consistency of variable dependencies. However, as the RCM’s output is physically consistent, similarity of the ML’s samples to the RCM implicitly evaluates this. Future work could focus more on physics-based evaluation.
@@ 6 Variability in EnScale’s samples
EnScale naturally allows us to separate sources of variability on different scales, as it splits the full GCM-RCM map into two steps. First, the coarse model maps GCM predictors to average-pooled RCM data and second, the super-resolution model learns the distribution of the high-resolution target given the average-pooled predictors (Fig. 2). This means we can split the full variability into contributions from the two paths: (i) coarse-scale variability, quantifying uncertainties in the pooled RCM data given GCM inputs, and (ii) local-scale variability, capturing variance in high-resolution RCM fields when conditioned on their coarse averages (see Sec. 4.3.1 for the formal variance decomposition).
For a first qualitative insight, we first compare EnScale’s full GCM-RCM map (as in the examples in Fig. 7) with examples from the pure super-resolution model. Using average-pooled versions of the target RCM data as inputs, EnScale’s samples agree very closely with the target (see Fig. A.11). The different samples show only small discrepancies on a local scale and variability is greatly reduced in this case compared to the full GCM-RCM model. This highlights that the full RCM emulation is a very different task from a pure super-resolution setup. In addition, it indicates that variability between GCM and RCM on coarse scales is a main source of the mismatches that we observed in samples by EnScale (see Fig. 7) and benchmarks (see Fig. A.6A.7).
The quantitative result of the decomposition is presented in Tab. 7, for each variable separately, averaged over locations. For all variables, the ratio of coarse-scale to local-scale variance greatly exceeds 1, indicating that most variability stems from the transfer of GCMs to average-pooled RCMs. In other words, the GCM boundary conditions contain comparatively little information about the coarsened RCM data, whereas the coarsened RCM already well determines the high-resolution output.
@@
Individual variances can be interpreted. We consider temperature, which is approximately normally distributed. Calculating $\pm 2\sqrt{\operatorname{Var}}$ gives the range around the mean in which approx. 95 % of data points lie. For example, for the local-scale variance $\pm 2\sqrt{\operatorname{Var}}=\pm 0.98$. That means, after fixing the average-pooled RCM data, local temperatures fluctuate by around $\pm$ $1\text{\,}\mathrm{K}$ in each grid point. Conditioning on the GCM input, however, temperatures on coarse scales can still vary by about $\pm$ $4.0\text{\,}\mathrm{K}$.
Local-scale variance in the RCM is to be expected, as the average values over larger areas can never determine the high-resolution output uniquely. Nevertheless, these findings indicate that variability on coarse scales is even more prominent. The GCM input gives only small constraints on the average-pooled RCM data. We suspect that this is because the GCM only drives the RCM at the boundaries, leaving a wide range of plausible climate conditions inside the RCM’s domain.
@@
The variances here can be interpreted as well, similar to the discussion in Sec. 6.1. For example, for temperature, the variance of one RCM is $3.9\text{\,}\mathrm{K}^{2}$. That means, for a fixed GCM input, temperatures simulated by the RCM in each location can fluctuate by $\pm$4.2\text{\,}\mathrm{K}$\$ (based on an approximation with a normal distribution, and calculated as above as $2\sqrt{\operatorname{Var}}$). However, the disagreements between RCMs are much smaller than that. In each location, the means of the RCMs differ by about $1.3\text{\,}\mathrm{K}$, when conditioning on a GCM input (calculated as $2\sqrt{\text{model diff.}}$, compare with the formula (5) in Sec. 4.3.2).
This finding points to the fact that internal variability seems to dominate over structural model differences for regional climate model ensembles. It is similar to [Rampal2025DownscalingExtremes], who use an RCM-emulator to present an uncertainty quantification into internal variability, scenario uncertainty and model uncertainty and also find a prominent role of internal variability. In addition, our variance decomposition underlines the benefits of jointly training one ML model for downscaling based on multiple GCM-RCM combinations. Even though inter-model differences can be rather large, intrinsic conditional variance of $Y|X$ dominates. Hence, there is sufficient similarity between GCM-RCM pairs such that the ML training can benefit from a joint training.

### v3

@@ 5.3.5 Extremes
Figure 9: Extreme quantiles for the summer season (June, July, August) compared to ALADIN63 driven by CNRM-CM5. The first row shows the 99 % - quantile of the RCM calculated separately in each location. The remaining rows show the error for different methods, calculated as the signed difference of quantiles in the samples minus the RCM’s quantiles. Additionally, the last row provides an assessment of “model uncertainty” where the difference of ALADIN63 to CCLM4-8-17 (both driven by CNRM-CM5) is shown. Similar results are obtained for pairs other than ALADIN63 driven by CNRM-CM5.
@@
We focus on the capability of EnScale and the benchmark to capture extreme events. As an example, we show the errors for the 99 % - quantile in summer months and for the first GCM-RCM pair in Fig. 9. The figure shows the true quantile in the RCM data (first row), which varies spatially, e.g. with largest extremes for precipitation and solar radiation over the Alps. Also, the magnitude and direction of errors vary across locations and variables. For example, EnScale overestimates extreme temperatures but underestimates extreme solar radiation in the majority of locations. For precipitation, absolute errors are highest in the Alpine region, and for surface wind, errors are largest over the Sea, where also the variability of the data is highest. Again, EnScale performs well in comparison to the benchmarks, with errors of similar magnitude as for the analogues and CorrDiff-s, and often lower than for GAN and CorrDiff.
We note, however, that error estimates are conservative – per season, there are approximately $1000$ days in the test set, leading to estimation errors for extreme quantiles. Thus, the error presented is a mixture of mistakes due to finite-sample approximation and actual limitation of the ML models and benchmarks.
@@
In order to provide an additional perspective on these results, we compare to uncertainties across different RCMs. As an example, we show differences between the quantiles of two RCMs, both driven by the same GCM. We find that the mismatches between RCMs are of similar magnitude to the errors of the ML models, suggesting that the emulators produce results falling into the range spanned by multiple RCMs.
In principle, if the ML models perfectly emulated each RCM individually, one would expect disagreements between RCMs to exceed the errors of the ML methods. This is mostly the case for EnScale, EnScale-t, CorrDiff and CorrDiff-s. However, the joint training across multiple GCM-RCM pairs likely also influences the learned distribution: Especially for extreme events, where there is only little data per individual RCM, our method seems to partly emulate the pooled distribution across multiple pairs instead of exactly reproducing the RCM-specific characteristics. This effect could also be related to our relatively simple approach for RCM conditioning (Sec. C.4) and could possibly be improved in future work.
Further evaluation is shown in Tab. A.2. We assess locationwise quantiles. EnScale performs well compared to the benchmarks, but best scores are often reached by the analogues and EasyUQ. Notably, the GAN and CorrDiff struggle to reproduce extreme quantiles accurately. In particular, they miss to predict the zero precipitation events correctly. The true 1 % -quantile in winter is zero, but the ML approaches do not capture this. EnScale successfully predicts dry days with zero rain. This is an advantage of our data transformation (Sec. 3.2).
@@
In addition, we consider calibration of low / high quantiles of the predicted conditional distribution (see “MCB Q” columns in Tab. A.2 and Sec. 4.2.4 for details). Here, EnScale, EnScale-t and CorrDiff-s usually reach the lowest errors, with the exact ordering of these three methods depending on the variable and quantile considered. This again exemplifies that EnScale is sometimes inferior to the analogues with respect to metrics evaluating the marginal distributions $p_{Y}$. For the conditional $p_{Y|X}$, however, EnScale is superior.
Evaluating EnScale for impact-relevant extremes leaves multiple possible directions of future work. Preliminary analysis suggests that EnScale can capture compound extremes [15]. We are aware that our choice of quantiles is only moderately extreme, and evaluating events with higher return periods would be relevant, but also require a longer test data set. Improving performance for extremes with, e.g., weighted proper scoring rules as loss functions is also a promising direction [75, 76].
@@ 5.4 Variable dependencies
So far, our evaluation considered each target variable separately. Next, we evaluate the capability of our model and the benchmarks to capture the full distribution jointly for the four targets in Tab. A.3. We find that EnScale and EnScale-t are often the best-performing ML methods, reaching similar scores as CorrDiff-s. More details can be found in App. A.5.
@@
We note that the above metrics only evaluate pairwise distributions between variables. Assessing the full joint distribution is not considered. In addition, we do not explicitly evaluate physical consistency of variable dependencies. However, as the RCM’s output is physically consistent, similarity of the ML’s samples to the RCM implicitly evaluates this. Future work could focus more on physics-based evaluation.
5.5 Temporal structure
Figure 10: Time series examples. Solid lines show the RCM time series (green) and a single randomly chosen realization of EnScale (blue), EnScale-t (yellow), CorrDiff (orange) and CorrDiff-s (light blue). Shaded areas indicate the minimum and maximum from 9 samples of each model. The figure shows spatially averaged temperature for the first year in the interpolation test set (2030) and for the first GCM-RCM pair (CNRM-CM5 and ALADIN63).
We investigate the temporal structure of the generated samples, that is, how accurately the methods capture correlation between consecutive time steps. Fig. 10 presents example time series for the spatial mean of temperature over one year. We compare EnScale, EnScale-t, CorrDiff, CorrDiff-s, and the underlying RCM.  [^4] The remaining variables and time series for EnScale, EnScale-t, and the underlying RCM are shown in Fig. A.13. Most importantly, we find that generated time series by both, EnScale and EnScale-t, closely match the ground truth over the entire year and capture the seasonal cycle. On most days, the truth lies within the range of the 9 samples.
The figure also highlights qualitative differences between generated time series by EnScale-t and EnScale. The time series for the EnScale-t tend to be smoother because they take into account the temporal structure, especially for temperature. The time series by EnScale is not necessarily always consistent over time as each day is generated independently of the previous. The uncertainty bands are similar for both approaches. In times where temporal autocorrelation is high (especially visible for temperature), the uncertainty range for the EnScale-t is larger because including temporal structure allows us to sample a larger variety of possible trajectories. Comparing CorrDiff and CorrDiff-s highlights the effect of the sampling scheme: The original implementation for CorrDiff generates smooth time series with large temporal correlations, while independent sampling in CorrDiff-s leads to more “noisy” time series, qualitatively similar to EnScale.
These results also show the stability of the time series of the autoregressive rollout for our temporally-consistent coarse model. Both the individual time series as well as the uncertainty range remain stable over time. We also tested longer rollout of our generated time series and found that they are stable even for several years (not shown). Both the time series as well as the loss do not increase over time. The stability of rollout is an open discussion in the case of ML-based weather forecasting [49, 32]. In our case, the additional input from the GCM likely helps to stabilize the rollout and prevent potential drifts.
Figure 11: Autocorrelation functions (ACFs). We calculate ACFs for lags up to 20 days for the raw time series (top row) and for a time series after high-pass filtering (bottom row), i.e. removing a 30-day-running mean from the time series before ACF calculation. ACFs are averaged across all locations and all GCM-RCM pairs.
Autocorrelation functions (ACFs) are shown in Fig. 11. EnScale underestimates local autocorrelation for all variables, with the effect being strongest for temperature. EnScale-t is able to reduce the error substantially and matches true ACFs very closely, both for the raw time series as well as after high-pass filtering.
To compare EnScale and EnScale-t to the benchmarks, Tab. A.4 shows both signed as well as absolute errors in lag-1 (i.e., a single time step) ACFs. We find that EnScale-t is superior to the others and achieves the lowest deviation from the true ACFs across all variables. All methods except NN-det and CorrDiff have negative differences, i.e., they underestimate the local autocorrelation. CorrDiff accurately reproduces autocorrelations for temperature, but overestimates them for all other variables.
Capturing autocorrelations is particularly relevant when analyzing statistics aggregated over multiple days, such as in 5-day heatwaves. Underestimating autocorrelation can lead to an underrepresentation of variability in multi-day averages, therefore potentially underestimating risk. These findings point to limitations of existing ML-based downscaling approaches that fail to include temporal correlation on a local scale.
5.6 Evaluation on extrapolation test period
Figure 12: Climate change trend for the average temperature between 2030-2039 and 2090-2099. The trend is computed per GCM-RCM pair first and then averaged across pairs.
So far, the evaluation has focused on the interpolation test period (2030-2039). We also provide scores on the extrapolation test set (2090-2099) (see Sec. A.7). We find that scores are very similar and the conclusions are consistent. This shows that our method is able to extrapolate in a mild extrapolation case. Note that training data extends until 2089, so 2090-99 represents only a small extrapolation. For example, in GCM-RCM pairs with relatively strong climate change signal, the extrapolation test data is, on average, only $0.77\text{\,}\mathrm{K}$ warmer than the last decade of the train data, with warming of around $1.2\text{\,}\mathrm{K}$ in a few locations.
We also show climate change patterns in Fig. 12. We present only EnScale; EnScale-t yields similar results. For the mean shift, errors of EnScale are small. We slightly underestimate the climate change signal over land, by around 5 or at most 10 % of the true climate change signal. Considering climate change trends in the extremes, we find that EnScale broadly captures the trend in the RCM. However, the spatial pattern of the temperature trend is not correct for the samples: Over land, the trend is underestimated.
We checked climate change trends also for other variables, but we found that there are large inter-model differences, yielding very small absolute values for the average climate change signals. This lack of clear signal in the data also leads to larger relative errors for EnScale, and we found it difficult to reach robust conclusions.
Our current data transformation (Sec. 3.2) relies on fitted ECDFs and samples data within the support of the training data. Thus, it is not designed for strong out-of-support-extrapolation setups. This might be one reason for the slight underestimation of climate change signals for temperature. Further improving the extrapolation capabilities of EnScale would require an adjustment of the data transformation, which we leave to future work.
5.7 Computational cost
Table 3: Training and inference times for EnScale, the GAN, and CorrDiff/CorrDiff-sa.
Method Training Inference
EnScale 26 h 1.75 h
EnScale-t 31.4 h 1.8 h
CorrDiff(-s) 216 h 36 h
GAN 125 h 20 h
a All runtimes were checked on a single NVIDIA A100 GPU (80GB) and exclude data loading. Inference times refer to producing a high-resolution ensemble of 100 years for a single GCM-RCM pair with 10 samples per day.
EnScale aims to be a computationally lightweight alternative to state-of-the-art generative models. As an example, we compare train and inference times of EnScale to the best-performing and most recent ML-baseline, CorrDiff (Tab. 3). CorrDiff-s only adjusts the sampling scheme compared to CorrDiff, so train and inference times are identical. We find that training EnScale only takes about 1 day, whereas the diffusion model requires almost 10 times as long. The difference for inference is even larger, with EnScale being around 20 times faster. However, it should be noted that inference times can likely be improved for both models. For EnScale, the sampling can be done in a few minutes, but the current bottleneck is the inversion of the data-transformation. EnScale-t is only slightly more computationally intensive than EnScale. Note that, in the current implementation, we also use EnScale to generate a starting value for EnScale-t, hence, sampling from EnScale-t requires training both EnScale and EnScale-t.
Train times depend on the batch size and stopping criterion. We train all models until convergence, i.e. until the test loss flattens. For CorrDiff, the deterministic UNet and the stochastic residual model both are trained with a batch size of 32. Training the residual model is more expensive and takes around 0.8s/batch or, equivalently, 38 samples/second. During training of the GAN, we use a batch size of 16, and the model takes around 3s/batch or 5 samples/second. For EnScale, the last super-resolution step is the most computationally intensive. It is trained with a batch size of 128, taking around 0.2s/batch or 581 samples/second.
EnScale is parameter-efficient due to its sparse layers (Sec. 2.4). Despite modeling effects specific to each location, the network only has 63 million parameters. EnScale-t needs an additional 3 million parameters, yielding 66 million in total. The UNet in our CorrDiff implementation has around 80 million parameters (as in [44]) and is used in both the regression as well as the residual modeling step, yielding a total of around 160 million parameters. The GAN has around 69 million parameters.
Computational efficiency gains for EnScale stem from both the overall setup and the model architecture. The sparse local layers were designed for progressive, stochastic resolution refinement, are effective when combined with the energy score loss and our multi-step setup (Sec. 2.2), and are key to achieving the short training time. For example, training a fully convolutional architecture for the last steps of the super-resolution model took several days.
@@ 6 Variability in EnScale’s samples
EnScale naturally allows us to separate sources of variability on different scales, as it splits the full GCM-RCM map into two steps. First, the coarse model maps GCM predictors to average-pooled RCM data and second, the super-resolution model learns the distribution of the high-resolution target given the average-pooled predictors (Fig. 2). This means we can split the full variability into contributions from the two parts: (i) coarse-scale variability, quantifying uncertainties in the pooled RCM data given GCM inputs, and (ii) local-scale variability, capturing variance in high-resolution RCM fields when conditioned on their coarse averages (see Sec. 4.3.1 for the formal variance decomposition).
For a first qualitative insight, we present samples of the pure EnScale’s super-resolution model, where average-pooled versions of the target RCM data are used as inputs and the coarse correction step is bypassed. In this setup, the samples agree very closely with the target (see Fig. A.10), and show much smaller variability than the samples from EnScale’s full GCM-RCM map (as shown in Fig. 5). The different samples from the super-resolution model show only small discrepancies on a local scale and variability is greatly reduced in this case compared to the full GCM-RCM model. This highlights that the full RCM emulation is a very different task from a pure super-resolution setup. In addition, it indicates that variability between GCM and RCM on coarse scales is a main source of the mismatches that we observed in samples by EnScale (see Fig. 5) and benchmarks (see Fig. A.3A.4).
The quantitative result of the decomposition is presented in Tab. 4, for each variable separately, averaged over locations. For all variables, the ratio of coarse-scale to local-scale variance greatly exceeds 1, indicating that most variability stems from the transfer of GCMs to average-pooled RCMs. In other words, the GCM boundary conditions contain comparatively little information about the coarsened RCM data, whereas the coarsened RCM already well determines the high-resolution output. We suspect that this is because the GCM only drives the RCM at the boundaries, leaving a wide range of plausible climate conditions inside the RCM’s domain.
@@
Individual variances can be interpreted. We consider temperature, which is approximately normally distributed. Calculating $\pm 2\sqrt{\operatorname{Var}}$ gives the range around the mean in which approx. 95 % of data points lie. For example, for the local-scale variance $\pm 2\sqrt{\operatorname{Var}}=\pm 0.98$. That means, after fixing the average-pooled RCM data, local temperatures fluctuate by around $\pm$ $1\text{\,}\mathrm{K}$ in each grid point. Conditioning on the GCM input, however, temperatures on coarse scales can still vary by about $\pm$ $3.7\text{\,}\mathrm{K}$.
@@
The variances here can be interpreted as well, similar to the discussion in Sec. 6.1. For example, for temperature, the variance of one RCM is $4.1\text{\,}\mathrm{K}^{2}$. That means, for a fixed GCM input, temperatures simulated by the RCM in each location can fluctuate by $\pm$4.0\text{\,}\mathrm{K}$\$ in an approximate 95 % confidence interval (based on an approximation with a normal distribution, and calculated as above as $2\sqrt{\operatorname{Var}}$). However, the disagreements between RCMs are much smaller than that. In each location, the means of the RCMs differ by about $1.3\text{\,}\mathrm{K}$, when conditioning on a GCM input (calculated as $2\sqrt{\text{model diff.}}$, compare with the formula (4) in Sec. 4.3.2).
This finding points to the fact that internal variability seems to dominate over structural model differences for regional climate model ensembles. It is similar to [50], who use an RCM-emulator to present an uncertainty quantification into internal variability, scenario uncertainty and model uncertainty and also find a prominent role of internal variability. In addition, our variance decomposition underlines the benefits of jointly training one ML model for downscaling based on multiple GCM-RCM combinations. Even though inter-model differences can be rather large, intrinsic conditional variance of $Y|X$ dominates. Hence, there is sufficient similarity between GCM-RCM pairs such that the ML training can benefit from a joint training.